# How Transformers Solve Propositional Logic Problems: A Mechanistic Analysis

### v3

@@ Appendix A Propositional logic problem and examples
Appendix B Length-3 small transformer study: experimental details
B.1 Data definition and examples
@@
Figure 6: Reasoning accuracies of several models on the length-3 problem. x-axis: model architecture (number of layers, number of heads); y-axis: reasoning accuracy. Note that the 3-layer 3-head variant is the smallest which obtains 100% accuracy on the logic problems.
B.2 Small-transformer characteristics, and training details
B.2.1 Transformer definition
@@
The following is the more technical definition of the model. Define input ${\bm{x}}=(x_{1},x_{2},...,x_{t})\in\mathbb{N}^{t}$, a sequence of tokens with length $t$. It is converted into a sequence of (trainable) token embeddings ${\bm{X}}_{token}=({\bm{e}}(x_{1}),{\bm{e}}(x_{2}),...,{\bm{e}}(x_{t}))^{T}\in\mathbb{R}^{t\times d_{e}}$, where we denote the hidden embedding dimension of the model with $d_{e}$. Adding to it the (trainable) positional embeddings ${\bm{P}}=({\bm{p}}_{1},{\bm{p}}_{2},...,{\bm{p}}_{t})^{T}\in\mathbb{R}^{t\times d_{e}}$, we form the zero-th layer embedding of the transformer
${\bm{X}}_{0}={\bm{X}}_{token}+{\bm{P}}=({\bm{e}}(x_{1})+{\bm{p}}_{1},...,{\bm{e}}(x_{t})+{\bm{p}}_{t}).$ (2)
@@
$\mathcal{A}_{\ell,j}({\bm{X}}_{\ell-1})=\mathcal{S}\left(\text{causal}\left[\frac{1}{\sqrt{d_{h}}}\left({\bm{Q}}_{\ell,j}\widetilde{{\bm{X}}}_{\ell-1}^{T}\right)^{T}{\bm{K}}_{\ell,j}\widetilde{{\bm{X}}}_{\ell-1}^{T}\right]\right)\widetilde{{\bm{X}}}_{\ell-1}{\bm{V}}_{\ell,j}^{T}\in\mathbb{R}^{t\times d_{h}},$ (3)
@@
  • $\widetilde{{\bm{X}}}_{\ell-1}=\text{LayerNorm}({\bm{X}}_{\ell-1})\in\mathbb{R}^{t\times d_{e}}$, where LayerNorm denotes the layer normalization operator (Ba et al., 2016).
  • ${\bm{Q}}_{\ell,j},{\bm{K}}_{\ell,j}\in\mathbb{R}^{d_{h}\times d_{e}}$ are the key and query matrices of attention head $(\ell,j)$, where $d_{h}=\frac{d_{e}}{H}$. They are multiplied with the input $\widetilde{{\bm{X}}}_{\ell-1}$ to obtain the query and key activations ${\bm{Q}}_{\ell,j}\widetilde{{\bm{X}}}_{\ell-1}^{T}$ and ${\bm{K}}_{\ell,j}\widetilde{{\bm{X}}}_{\ell-1}$, both in the space $\mathbb{R}^{d_{h}\times t}$. We then perform the “scaled dot-product” of the query and key activations to obtain
    $\frac{1}{\sqrt{d_{h}}}\left({\bm{Q}}_{\ell,j}\widetilde{{\bm{X}}}_{\ell-1}^{T}\right)^{T}{\bm{K}}_{\ell,j},\widetilde{{\bm{X}}}_{\ell-1}^{T}\in\mathbb{R}^{t\times t},$ (4)
@@
  • The causal mask operator $\text{causal}:\mathbb{R}^{t\times t}\to\mathbb{R}^{t\times t}$ allows the lower triangular portion of the input (including the diagonal entries) to pass through unchanged, and sets the upper triangular portion of the input to $-U$, where $U$ is a very large positive number (some papers simply denote this $-U$ as $\infty$- ∞). In other words, given any ${\bm{M}}\in\mathbb{R}^{t\times t}$ and $(i,k)\in[t]\times[t]$,
    $\left[\text{causal}\left[{\bm{M}}\right]\right]_{i,k}=\left[{\bm{M}}\right]_{i,k},\,\text{if }i\geq k;$ (5)
    $\left[\text{causal}\left[{\bm{M}}\right]\right]_{i,k}=-U,\,\text{if }i<k.$
  • $\mathcal{S}:\mathbb{R}^{t\times t}\to\mathbb{R}^{t\times t}$ is the softmax operator, which computes the row-wise softmax output from the input square matrix. In particular, given a square input matrix ${\bm{M}}\in\mathbb{R}^{t\times t}$ with its upper triangular portion set to $-U$ (note that the causal mask operator indeed causes the input to $\mathcal{S}$ to have this property), we have
    $[\mathcal{S}({\bm{M}})]_{i,k}=\frac{\exp\left([{\bm{M}}]_{i,k}\right)}{\sum_{n=1}^{i}\exp([{\bm{M}}]_{i,n})},\,\text{if }i\geq k;$ (6)
    $[\mathcal{S}({\bm{M}})]_{i,k}=0,\,\text{if }i<k.$
  • To recap a bit, we have now explained how to compute the first major term in equation 3, namely $\mathcal{S}\left(\text{causal}\left[\frac{1}{\sqrt{d_{h}}}\left({\bm{Q}}_{\ell,j}\widetilde{{\bm{X}}}_{\ell-1}^{T}\right)^{T}{\bm{K}}_{\ell,j}\widetilde{{\bm{X}}}_{\ell-1}^{T}\right]\right)\in[0,1]^{t\times t}$. It reflects the attention pattern (also called attention probabilities) of the attention head $(\ell,j)$ illustrated in Figure 7’s right half. Intuitively speaking, the $(i,k)$ entry of this $t$ by $t$ matrix reflects how much the attention head moves the information from the previous layer $\ell-1$ at the source token position of $k$ to the current layer $\ell$ at the target token position $i$.
  • Now what about $\widetilde{{\bm{X}}}_{\ell-1}{\bm{V}}_{\ell,j}^{T}$? ${\bm{V}}_{\ell,j}\in\mathbb{R}^{d_{h}\times d_{e}}$ is the value matrix of attention head $(\ell,j)$. It is multiplied with $\widetilde{{\bm{X}}}_{\ell-1}$ to obtain the value activation $\widetilde{{\bm{X}}}_{\ell-1}{\bm{V}}_{\ell,j}^{T}\in\mathbb{R}^{t\times d_{h}}$.
  • At this point, we have shown how the whole term in equation equation 3 is computed.
@@
${\bm{X}}_{\ell}={\bm{X}}_{\ell-1}+\text{Concat}[\mathcal{A}_{\ell,1}({\bm{X}}_{\ell-1}),...,\mathcal{A}_{\ell,H}({\bm{X}}_{\ell-1})]{\bm{W}}_{O,\ell}^{T}.$ (7)
@@
  • $\text{Concat}[\cdot]$ is the concatenation operator, where $\text{Concat}[\mathcal{A}_{\ell,1}({\bm{X}}_{\ell-1}),...,\mathcal{A}_{\ell,H}({\bm{X}}_{\ell-1})]\in\mathbb{R}^{t\times d_{e}}$.
  • ${\bm{W}}_{O,\ell}\in\mathbb{R}^{d_{e}\times d_{e}}$ is the projection matrix (sometimes called output matrix) of layer $\ell$. In our implementation, we allow this layer to have trainable bias terms too.
Finally, having computed, layer by layer, the hidden outputs ${\bm{X}}_{\ell,t}$ for $\ell\in[L]$, we apply an affine classifier (with softmax) to obtain the output of the model
${\bm{f}}({\bm{x}})=\mathcal{S}(\widetilde{{\bm{X}}}_{L,t}{\bm{W}}_{class}^{T}+{\bm{b}}_{class})$ (8)
@@ B.2.2 Training details
B.3 High-level reasoning strategy of the 3-layer transformer
@@
Figure 8: High-level overview of how the 3-layer transformer solves the logic problem. As shown in the grey blocks on the left, the model performs “routing” in layer 2 by sending a routing signal ${\bm{h}}_{route}$ to layer 3 (with its “sign” dependent on the query type), then the layer-3 attention block acts according to the “sign” of the routing signal sent to it. The middle (right) chain shows the strategy when the problem queries for linear (LogOp) chain.
B.4 Answer for the LogOp Chain
Evidence 3a: Distinct behaviors of affine predictors at different layers. We train two affine classifiers at two positions inside the model (each with 10k samples): ${\bm{W}}_{resid,\ell=2}$ at layer-2 residual stream, and ${\bm{W}}_{attn,\ell=3}$ at layer-3 attention-block output, both at the position of ANSWER, with the target being the correct first token. In training, if there are two correct answers possible (e.g. OR gate, starting nodes are both TRUE or both FALSE), we randomly choose one as the target; in testing, we deem the top-1 prediction “correct” if it coincides with one of the answers. We observe the following predictor behavior on the test samples:
@@ B.4 Answer for the LogOp Chain
    ${\bm{W}}_{attn,\ell=3}$ predicts the correct answer 100% of the time.
@@
    ${\bm{W}}_{resid,\ell=2}$ always predicts one of the variables assigned FALSE (in the fact section) if LogOp is the AND gate, and predicts one assigned TRUE if LogOp is the OR gate.
@@
  • Construct two sets of samples $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, each of size 10k: for every sample ${\bm{X}}_{1,n}\in\mathcal{D}_{1}$ and ${\bm{X}}_{2,n}\in\mathcal{D}_{2}$, the context of the two samples are exactly the same, except the LogOp is flipped, i.e. if ${\bm{X}}_{1,n}$ has disjunction, then ${\bm{X}}_{2,n}$ has the conjunction operator. If layer 3 of the model has no reliance on the $\text{Resid}_{\ell=2}$ (layer-2 residual stream) for LogOp information at the ANSWER position, then when we run the model on any ${\bm{X}}_{2,n}$, patching $\text{Resid}_{\ell=2}({\bm{X}}_{n,2})$ with $\text{Resid}_{\ell=2}({\bm{X}}_{n,1})$ at ANSWER should not cause significant change to the model’s accuracy of prediction. However, we observe the contrary: the accuracy of prediction degrades from 100% to 70.87%, with standard deviation 3.91% (repeated over 3 sets of experiments).
Observation: LogOp-relevant reasoning at the third layer. We show that the output from attention heads (3,1) and (3,3) (before the output/projection matrix of the layer-3 attention block), namely $\mathcal{A}_{3,1}({\bm{X}}_{2})$ and $\mathcal{A}_{3,3}({\bm{X}}_{2})$, when concatenated, contain linearly decodable information about the two starting nodes of the LogOp chain. We frame this as a multi-label classification problem, detailed as follows:
@@
    We train an affine classifier, with its input being the concatenated $\text{Concat}[\mathcal{A}_{3,1}({\bm{X}}_{2}),\mathcal{A}_{3,3}({\bm{X}}_{2})]$ (a 512-dimensional vector) on every training sample, and with the targets and training loss defined above. We use a constant learning rate of $0.5\times 10^{-3}$, and weight decay of $10^{-2}$. The optimizer is AdamW in PyTorch.

### v4

@@ C.2.2 Answer for the linear and LogOp chain
Appendix D Length-3 small transformer study: further technical details
In this section, we provide further technical details of our three-layer transformer experiments.
D.1 Data definition and examples
@@
Figure 27: Small transformers (trained from scratch), varying sizes. Reasoning accuracies of several models on the length-3 problem. x-axis: model architecture (number of layers, number of heads); y-axis: reasoning accuracy. Note that the 3-layer 3-head variant is the smallest which obtains 100% accuracy on the logic problems.
D.2 Small transformer characteristics, and training details
D.2.1 Transformer definition
@@
The following is the more technical definition of the model. Define input $\boldsymbol{x}=(x_{1},x_{2},...,x_{t})\in\mathbb{N}^{t}$, a sequence of tokens with length $t$. It is converted into a sequence of (trainable) token embeddings $\boldsymbol{X}_{token}=(\boldsymbol{e}(x_{1}),\boldsymbol{e}(x_{2}),...,\boldsymbol{e}(x_{t}))^{T}\in\mathbb{R}^{t\times d_{e}}$, where we denote the hidden embedding dimension of the model with $d_{e}$. Adding to it the (trainable) positional embeddings $\boldsymbol{P}=(\boldsymbol{p}_{1},\boldsymbol{p}_{2},...,\boldsymbol{p}_{t})^{T}\in\mathbb{R}^{t\times d_{e}}$, we form the zero-th layer embedding of the transformer
$\boldsymbol{X}_{0}=\boldsymbol{X}_{token}+\boldsymbol{P}=(\boldsymbol{e}(x_{1})+\boldsymbol{p}_{1},...,\boldsymbol{e}(x_{t})+\boldsymbol{p}_{t}).$ (2)
@@
$\mathcal{A}_{\ell,j}(\boldsymbol{X}_{\ell-1})=\mathcal{S}\left(\text{causal}\left[\frac{1}{\sqrt{d_{h}}}\left(\boldsymbol{Q}_{\ell,j}\widetilde{\boldsymbol{X}}_{\ell-1}^{T}\right)^{T}\boldsymbol{K}_{\ell,j}\widetilde{\boldsymbol{X}}_{\ell-1}^{T}\right]\right)\widetilde{\boldsymbol{X}}_{\ell-1}\boldsymbol{V}_{\ell,j}^{T}\in\mathbb{R}^{t\times d_{h}},$ (3)
@@
  • $\widetilde{\boldsymbol{X}}_{\ell-1}=\text{LayerNorm}(\boldsymbol{X}_{\ell-1})\in\mathbb{R}^{t\times d_{e}}$, where LayerNorm denotes the layer normalization operator (Ba et al., 2016).
  • $\boldsymbol{Q}_{\ell,j},\boldsymbol{K}_{\ell,j}\in\mathbb{R}^{d_{h}\times d_{e}}$ are the key and query matrices of attention head $(\ell,j)$, where $d_{h}=\frac{d_{e}}{H}$. They are multiplied with the input $\widetilde{\boldsymbol{X}}_{\ell-1}$ to obtain the query and key activations $\boldsymbol{Q}_{\ell,j}\widetilde{\boldsymbol{X}}_{\ell-1}^{T}$ and $\boldsymbol{K}_{\ell,j}\widetilde{\boldsymbol{X}}_{\ell-1}$, both in the space $\mathbb{R}^{d_{h}\times t}$. We then perform the “scaled dot-product” of the query and key activations to obtain
    $\frac{1}{\sqrt{d_{h}}}\left(\boldsymbol{Q}_{\ell,j}\widetilde{\boldsymbol{X}}_{\ell-1}^{T}\right)^{T}\boldsymbol{K}_{\ell,j},\widetilde{\boldsymbol{X}}_{\ell-1}^{T}\in\mathbb{R}^{t\times t},$ (4)
@@
  • The causal mask operator $\text{causal}:\mathbb{R}^{t\times t}\to\mathbb{R}^{t\times t}$ allows the lower triangular portion of the input (including the diagonal entries) to pass through unchanged, and sets the upper triangular portion of the input to $-U$, where $U$ is a very large positive number (some papers simply denote this $-U$ as $\infty$- ∞). In other words, given any $\boldsymbol{M}\in\mathbb{R}^{t\times t}$ and $(i,k)\in[t]\times[t]$,
    $\left[\text{causal}\left[\boldsymbol{M}\right]\right]_{i,k}=\left[\boldsymbol{M}\right]_{i,k},\,\text{if }i\geq k;$ (5)
    $\left[\text{causal}\left[\boldsymbol{M}\right]\right]_{i,k}=-U,\,\text{if }i<k.$
  • $\mathcal{S}:\mathbb{R}^{t\times t}\to\mathbb{R}^{t\times t}$ is the softmax operator, which computes the row-wise softmax output from the input square matrix. In particular, given a square input matrix $\boldsymbol{M}\in\mathbb{R}^{t\times t}$ with its upper triangular portion set to $-U$ (note that the causal mask operator indeed causes the input to $\mathcal{S}$ to have this property), we have
    $[\mathcal{S}(\boldsymbol{M})]_{i,k}=\frac{\exp\left([\boldsymbol{M}]_{i,k}\right)}{\sum_{n=1}^{i}\exp([\boldsymbol{M}]_{i,n})},\,\text{if }i\geq k;$ (6)
    $[\mathcal{S}(\boldsymbol{M})]_{i,k}=0,\,\text{if }i<k.$
  • To recap a bit, we have now explained how to compute the first major term in (3), namely $\mathcal{S}\left(\text{causal}\left[\frac{1}{\sqrt{d_{h}}}\left(\boldsymbol{Q}_{\ell,j}\widetilde{\boldsymbol{X}}_{\ell-1}^{T}\right)^{T}\boldsymbol{K}_{\ell,j}\widetilde{\boldsymbol{X}}_{\ell-1}^{T}\right]\right)\in[0,1]^{t\times t}$. It reflects the attention pattern (also called attention probabilities) of the attention head $(\ell,j)$ illustrated in Figure 28’s right half. Intuitively speaking, the $(i,k)$ entry of this $t$ by $t$ matrix reflects how much the attention head moves the information from the previous layer $\ell-1$ at the source token position of $k$ to the current layer $\ell$ at the target token position $i$.
  • Now what about $\widetilde{\boldsymbol{X}}_{\ell-1}\boldsymbol{V}_{\ell,j}^{T}$? $\boldsymbol{V}_{\ell,j}\in\mathbb{R}^{d_{h}\times d_{e}}$ is the value matrix of attention head $(\ell,j)$. It is multiplied with $\widetilde{\boldsymbol{X}}_{\ell-1}$ to obtain the value activation $\widetilde{\boldsymbol{X}}_{\ell-1}\boldsymbol{V}_{\ell,j}^{T}\in\mathbb{R}^{t\times d_{h}}$.
  • At this point, we have shown how the whole term in equation (3) is computed.
@@
$\boldsymbol{X}_{\ell}=\boldsymbol{X}_{\ell-1}+\text{Concat}[\mathcal{A}_{\ell,1}(\boldsymbol{X}_{\ell-1}),...,\mathcal{A}_{\ell,H}(\boldsymbol{X}_{\ell-1})]\boldsymbol{W}_{O,\ell}^{T}.$ (7)
@@
  • $\text{Concat}[\cdot]$ is the concatenation operator, where $\text{Concat}[\mathcal{A}_{\ell,1}(\boldsymbol{X}_{\ell-1}),...,\mathcal{A}_{\ell,H}(\boldsymbol{X}_{\ell-1})]\in\mathbb{R}^{t\times d_{e}}$.
  • $\boldsymbol{W}_{O,\ell}\in\mathbb{R}^{d_{e}\times d_{e}}$ is the projection matrix (sometimes called output matrix) of layer $\ell$. In our implementation, we allow this layer to have trainable bias terms too.
Finally, having computed, layer by layer, the hidden outputs $\boldsymbol{X}_{\ell,t}$ for $\ell\in[L]$, we apply an affine classifier (with softmax) to obtain the output of the model
$\boldsymbol{f}(\boldsymbol{x})=\mathcal{S}(\widetilde{\boldsymbol{X}}_{L,t}\boldsymbol{W}_{class}^{T}+\boldsymbol{b}_{class})$ (8)
@@ D.2.2 Training details
D.3 High-level reasoning strategy of the 3-layer transformer
@@
Figure 29: High-level overview of how the 3-layer transformer solves the logic problem. As shown in the grey blocks on the left, the model performs “routing” in layer 2 by sending a routing signal $\boldsymbol{h}_{route}$ to layer 3 (with its “sign” dependent on the query type), then the layer-3 attention block acts according to the “sign” of the routing signal sent to it. The middle (right) chain shows the strategy when the problem queries for linear (LogOp) chain.
D.4 Answer for the LogOp Chain
Evidence 3a: Distinct behaviors of affine predictors at different layers. We train two affine classifiers at two positions inside the model (each with 10k samples): $\boldsymbol{W}_{resid,\ell=2}$ at layer-2 residual stream, and $\boldsymbol{W}_{attn,\ell=3}$ at layer-3 attention-block output, both at the position of ANSWER, with the target being the correct first token. In training, if there are two correct answers possible (e.g. OR gate, starting nodes are both TRUE or both FALSE), we randomly choose one as the target; in testing, we deem the top-1 prediction “correct” if it coincides with one of the answers. We observe the following predictor behavior on the test samples:
@@ D.4 Answer for the LogOp Chain
    $\boldsymbol{W}_{attn,\ell=3}$ predicts the correct answer 100% of the time.
@@
    $\boldsymbol{W}_{resid,\ell=2}$ always predicts one of the variables assigned FALSE (in the fact section) if LogOp is the AND gate, and predicts one assigned TRUE if LogOp is the OR gate.
@@
  • Construct two sets of samples $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, each of size 10k: for every sample $\boldsymbol{X}_{1,n}\in\mathcal{D}_{1}$ and $\boldsymbol{X}_{2,n}\in\mathcal{D}_{2}$, the context of the two samples are exactly the same, except the LogOp is flipped, i.e. if $\boldsymbol{X}_{1,n}$ has disjunction, then $\boldsymbol{X}_{2,n}$ has the conjunction operator. If layer 3 of the model has no reliance on the $\text{Resid}_{\ell=2}$ (layer-2 residual stream) for LogOp information at the ANSWER position, then when we run the model on any $\boldsymbol{X}_{2,n}$, patching $\text{Resid}_{\ell=2}(\boldsymbol{X}_{n,2})$ with $\text{Resid}_{\ell=2}(\boldsymbol{X}_{n,1})$ at ANSWER should not cause significant change to the model’s accuracy of prediction. However, we observe the contrary: the accuracy of prediction degrades from 100% to 70.87%, with standard deviation 3.91% (repeated over 3 sets of experiments).
Observation: LogOp-relevant reasoning at the third layer. We show that the output from attention heads (3,1) and (3,3) (before the output/projection matrix of the layer-3 attention block), namely $\mathcal{A}_{3,1}(\boldsymbol{X}_{2})$ and $\mathcal{A}_{3,3}(\boldsymbol{X}_{2})$, when concatenated, contain linearly decodable information about the two starting nodes of the LogOp chain. We frame this as a multi-label classification problem, detailed as follows:
@@
    We train an affine classifier, with its input being the concatenated $\text{Concat}[\mathcal{A}_{3,1}(\boldsymbol{X}_{2}),\mathcal{A}_{3,3}(\boldsymbol{X}_{2})]$ (a 512-dimensional vector) on every training sample, and with the targets and training loss defined above. We use a constant learning rate of $0.5\times 10^{-3}$, and weight decay of $10^{-2}$. The optimizer is AdamW in PyTorch.